OpenStudy (anonymous):

i had similar question on knot, the system is balanced so that the 67N block on the left and the 14kg block on the right dont quite slip. the suspended block has mass of 2kg and the angle is 44, Find the coefficient of friction u1 for the 2kg block for this to be true, find the coefficient of friction u2 for 14kg block for this to be true

OpenStudy (anonymous):

i use T-M2g=0 T=2*99.8),=19.6 Tx-(u1*n)=o M1g=n T=u1*n 19.6=u1*67 u1=.293